OpenStudy (anonymous):

\[Let\ \ f(x)=x^7+x^6+x^5+x^4+x^3+x^2+x+1\]\[Find\ the\ remainder\ when\ f(x^{16})\ is\ divided\ by\ f(x).\]

OpenStudy (experimentx):

\[ x^{7}+x^{6}+x^5+x^4+x^3+x^2+x+1 = \frac{x^8 - 1}{x - 1}\] \[ f(x^{16})(x-1)/(x^8 - 1)\]

OpenStudy (experimentx):

Wild Guess \[ 8(x-1)\] ???

OpenStudy (anonymous):

no

OpenStudy (experimentx):

lol ...

OpenStudy (anonymous):

The remainder is 8.

OpenStudy (anonymous):

yeah! :)

OpenStudy (experimentx):

Haha .. how??

OpenStudy (anonymous):

One way is to do very long division.

OpenStudy (anonymous):

It has an easy way to find the remainder.

OpenStudy (anonymous):

Stay tuned. I am trying to find it.

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=PolynomialRemainder [(x^(7*16)%2Bx^(6*16)%2Bx^(5*16)%2Bx^(4*16)%2Bx^(3*16)%2Bx^(2*16)%2Bx^(16)%2B1)%2C+(x^8-1)%2C+x]

OpenStudy (experimentx):

For some weird reason, the remainder from these two divisions are same http://www.wolframalpha.com/input/?i=PolynomialRemainder [%28x^%287*16%29%2Bx^%286*16%29%2Bx^%285*16%29%2Bx^%284*16%29%2Bx^%283*16%29%2Bx^%282*16%29%2Bx^%2816%29%2B1%29%2C+x^7%2Bx^6%2Bx^5%2Bx^4%2Bx^3%2Bx^2%2Bx%2B1%2C+x]

OpenStudy (experimentx):

oh ..

OpenStudy (anonymous):

It seems that is a special case of this general case: For any integer \( n \ge 2 \) if we put \[ f(x) =\sum_{i=0}^{n-1} x^i \] then \( f(x) \) divides \( f\left( x^{2n}\right)- n \)

OpenStudy (anonymous):

I still do not know how to prove it. I will post it as a new problem and see if someone can come up with a clever solution.